# Path-dependent Hamilton–Jacobi equations with $u$-dependence and time-measurable Hamiltonians

### v1

@@ 5. Comparison principle
(5.1) $\begin{split}&\underline{\mathcal{A}}^{L}w(s_{0},x_{0},\tilde{x}_{0}):=\{\varphi\in C^{1,1,1}([s_{0},T]\times\Omega\times\Omega):\,\exists T_{0}\in(t_{0},T]:\\ &\qquad 0=(\varphi-w)(s_{0},x_{0},\tilde{x}_{0})=\inf\{(\varphi-w)(t,x,\tilde{x}):\\ &\qquad\qquad\qquad(t,x,\tilde{x})\in[s_{0},T_{0}]\in\mathcal{X}^{L}(s_{0},x_{0})\times\mathcal{X}^{L}(s_{0},\tilde{x}_{0})\}\}.\end{split}$
@@
on $[0,T)\times\Omega\times\Omega$ with parameter $\Upsilon$ if $w$ is u.s.c. and, for every $(s,x,\tilde{x})\in[0,T)\times\Omega\times\Omega$ and every test function $\varphi\in\underline{\mathcal{A}}^{L}w(s,x,\tilde{x})$, we have
@@
Let $L\geq 0$, $u$ be a minimax $L$-subsolution and $v$ be a minimax $L$-supersolution of (4.1). Then $(t,x,\tilde{x})\mapsto w(t,x,\tilde{x}):=u(t,x)-v(t,\tilde{x})$, $[0,T]\times\Omega\times\Omega\to{\mathbb{R}}$, is a viscosity $L$-subsolution of (5.2) with parameter $\Upsilon=u$.
@@
Let $(s_{0},x_{0},\tilde{x}_{0})\in[0,T)\times\Omega\times\Omega$, $w(s_{0},x_{0},\tilde{x}_{0})>0$, $\varphi\in\underline{\mathcal{A}}^{L}w(s_{0},x_{0},\tilde{x}_{0})$, and $T_{0}\in(s_{0},T]$ such that (5.1) holds. Let $y_{0}:=u(s_{0},x_{0})$, $\tilde{y}_{0}:=v(s_{0},x_{0})$, $z:=\partial_{x}\varphi(s_{0},x_{0},\tilde{x}_{0})$, and $\tilde{z}:=-\partial_{\tilde{x}}\varphi(s_{0},x_{0},\tilde{x}_{0})$. By the minimax semisolution properties of $u$ and $v$, there exist $(x,y)\in\mathcal{Y}^{L}(s_{0},x_{0},y_{0},z)$ and $(\tilde{x},\tilde{y})\in\mathcal{Y}^{L}(s_{0},\tilde{x}_{0},\tilde{y}_{0},\tilde{z})$, such that, for every $t\in[s_{0},T]$,
(5.4) $\begin{split}&[u(t,x)-y_{0}]-[v(t,\tilde{x})-\tilde{y}_{0}]\geq[y(t)-y_{0}]-[\tilde{y}(t)-\tilde{y}_{0}]\\ &\qquad=\int_{s_{0}}^{t}\left[(x^{\prime}(s),z)-H(s,x,y(s),z)-(\tilde{x}^{\prime}(s),\tilde{z})+H(s,\tilde{x},\tilde{y}(s),\tilde{z})\right]\,{\mathrm{d}}s.\end{split}$
@@
$\Psi(s,x,\tilde{x}):=\begin{cases}\frac{\sup\limits_{t\leq s}\left|{x(t)-\tilde{x}(t)}\right|^{2}-\left|{x(s)-\tilde{x}(s)}\right|^{2}}{\sup\limits_{t\leq s}\left|{x(t)-\tilde{x}(t)}\right|^{2}}+\left|{x(s)-\tilde{x}(s)}\right|^{2}&\text{if $\sup\limits_{t\leq s}\left|{x(t)-\tilde{x}(t)}\right|>0$,}\\ 0&\text{otherwise.}\end{cases}$
This functional has been introduced in [zhou2020viscosity] (see also [GLP21AMO] for further details). By the arguments in [GLP21AMO, Appendix B], one can deduce that $\Psi\in C^{1,1,1}([s_{0},T]\times\Omega\times\Omega)$ with derivatives $\partial_{t}\Psi(s,x,\tilde{x})=0$ and
$\partial_{x}\Psi(s,x,\tilde{x})=-\partial_{\tilde{x}}\Psi(s,x,\tilde{x})$
=$=\begin{cases}\left(2-\frac{4\sup\limits_{t\leq s}\left|{x(t)-\tilde{x}(t)}\right|^{2}-\left|{x(s)-\tilde{x}(s)}\right|^{2}}{\sup\limits_{t\leq s}\left|{x(t)-\tilde{x}(t)}\right|^{2}}\right)\cdot[x(s)-\tilde{x}(s)]&\text{if $\sup\limits_{t\leq s}\left|{x(t)-\tilde{x}(t)}\right|>0$,}\\ 0&\text{otherwise.}\end{cases}$
@@
$\Phi_{\varepsilon}(s,x,\tilde{x}):=w(s,x,\tilde{x})-\frac{T-s}{T-s_{0}}\cdot\frac{M_{0}}{2}-\frac{1}{{\varepsilon}}\Psi(s,x,\tilde{x}).$
Fix a point $k_{\varepsilon}=(s_{\varepsilon},x_{\varepsilon},\tilde{x}_{\varepsilon})$ at which the u.s.c. map $\Phi_{\varepsilon}$ attains a maximum on the compact set $K:=[s_{0},T]\times\mathcal{X}^{L}(s_{0},x_{0})\times\mathcal{X}^{L}(s_{0},x_{0})$. Note that
$M_{\varepsilon}:=\Phi_{\varepsilon}(k_{\varepsilon})\geq\Phi_{\varepsilon}(s_{0},x_{0},x_{0})=w(s_{0},x_{0},x_{0})-\frac{M_{0}}{2}=\frac{M_{0}}{2}.$
@@
(5.6) $w(k_{\varepsilon})\geq\frac{M_{0}}{2}+\frac{T-s_{\varepsilon}}{T-s_{0}}\cdot\frac{M_{0}}{2}+\frac{1}{{\varepsilon}}\Psi(k_{\varepsilon})\geq\frac{M_{0}}{2}>0.$
Moreover, $s_{\varepsilon}<T$ if ${\varepsilon}$ is sufficiently small (cf. Proposition 3.7 and its proof in [CrandallIshiiLions]).
Define a map $\varphi_{\varepsilon}:[s_{\varepsilon},T]\times\Omega\times\Omega\to{\mathbb{R}}$ by
$\varphi_{\varepsilon}(s,x,\tilde{x}):=M_{\varepsilon}+\frac{T-s}{T-s_{0}}\cdot\frac{M_{0}}{2}+\frac{1}{{\varepsilon}}\Psi(s,x,\tilde{x}).$
Then $\varphi_{\varepsilon}\in\underline{\mathcal{A}}^{L}w(k_{\varepsilon})$ with corresponding time $T_{0}=T$ because
$\varphi_{\varepsilon}(k_{\varepsilon})-w(k_{\varepsilon})=\overbrace{w(k_{\varepsilon})-\frac{T-s_{\varepsilon}}{T-s_{0}}\cdot\frac{M_{0}}{2}-\frac{1}{{\varepsilon}}\Psi(k_{\varepsilon})}^{=M_{\varepsilon}}$
$\qquad\qquad\qquad\qquad\qquad\qquad+\frac{T-s_{\varepsilon}}{T-s_{0}}\cdot\frac{M_{0}}{2}+\frac{1}{{\varepsilon}}\Psi(k_{\varepsilon})-w(k_{\varepsilon})=0$
$\qquad\leq M_{\varepsilon}-\Phi_{\varepsilon}(s,x,\tilde{x})=M_{\varepsilon}+\frac{T-s}{T-s_{0}}\cdot\frac{M_{0}}{2}+\frac{1}{{\varepsilon}}\Psi(s,x,\tilde{x})-w(s,x,\tilde{x})$
$\qquad=\varphi_{\varepsilon}(s,x,\tilde{x})-w(s,x,\tilde{x})$
@@
$\partial_{t}\varphi_{\varepsilon}(s,x,\tilde{x})=-\frac{M_{0}}{2(T-s_{0})}\quad\text{and}\quad\partial_{x}\varphi_{\varepsilon}(s,x,\tilde{x})=-\partial_{\tilde{x}}\varphi_{\varepsilon}(s,x,\tilde{x})=\frac{1}{{\varepsilon}}\partial_{x}\Psi(s,x,\tilde{x}).$
Consequently, since $w$ is a viscosity $L$-subsolution of (5.2) and since (5.6) holds,
(5.7) $\begin{split}0&\leq-\frac{M_{0}}{2(T-s_{0})}+M_{L}\left(1+\left|{\Upsilon(s_{\varepsilon},x_{\varepsilon})}\right|+\frac{1}{{\varepsilon}}\left|{\partial_{x}\Psi(k_{\varepsilon})}\right|\right)\cdot\sup_{t\leq s_{\varepsilon}}\left|{x_{\varepsilon}(t)-\tilde{x}_{\varepsilon}(t)}\right|\\ &\leq-\frac{M_{0}}{2(T-s_{0})}+\tilde{M}_{L}\left(1+\frac{1}{{\varepsilon}}\left|{\partial_{x}\Psi(k_{\varepsilon})}\right|\right)\cdot\sup_{t\leq s_{\varepsilon}}\left|{x_{\varepsilon}(t)-\tilde{x}_{\varepsilon}(t)}\right|\end{split}$
for some constant $\tilde{M}_{L}>0$ independent from ${\varepsilon}$. The second line of (5.7) follows from the upper semi-continuity of $\Upsilon$ and the compactness of $\mathcal{X}^{L}(s_{0},x_{0})$. It remains to note that (cf. [GLP21AMO, page S1102])
@@
for every $(s,x,\tilde{x})\in[0,T]\times\Omega\times\Omega$. Then we can deduce, as in the proof of Theorem 4.2 in [BK18JFA], that
$\left(1+\frac{1}{{\varepsilon}}\left|{\partial_{x}\Psi(k_{\varepsilon})}\right|\right)\cdot\sup_{t\leq s_{\varepsilon}}\left|{x_{\varepsilon}(t)-\tilde{x}_{\varepsilon}(t)}\right|\leq\sqrt{\Psi(k_{\varepsilon})}+\frac{C}{{\varepsilon}}\Psi(k_{\varepsilon})\to 0\quad\text{as ${\varepsilon}\downarrow 0$}$
@@
Let $L\geq 0$, $u$ be a minimax $L$-subsolution and $v$ be a minimax $L$-supersolution of (4.1). Then $u\leq v$ on $[0,T]\times\Omega$.
@@ 6. Existence
We show existence of minimax solutions via Perron’s method (see [BardiCapuzzoDolcetta, section V.2] regarding a corresponding treatment for non-continuous viscosity solutions in the non-path-dependent case). More precisely, the scheme888 This scheme, which provides existence for minimax solutions, actually predates “Perron’s method” for viscosity solutions introduced in [Ishii87Perron] (we refer to the discussion in [Subbotin, section 10.8] for more details). in [Subbotin, section 8] is adapted to path-dependent Hamilton-Jacobi equations. In the case of Hamiltonians without $u$-dependence, this has already been done in [Lukoyanov03, section 7] (cf. also [BK18JFA, section 5], whose structure we follow here).
@@
where $O_{\delta}(t_{0},x_{0})=\{(t,x)\in[0,T]\times\Omega:\,\mathbf{d}_{\infty}((t_{0},x_{0}),(t_{n},x_{n}))<\delta\}$. We show that $(u_{0})_{\ast}$ is a minimax $L_{H}$-supersolution of (4.1) and $(u_{0})^{\ast}$ is a minimax $L_{H}$-subsolution of (4.1). The comparison principle (Corollary 5.4) yields $(u_{0})^{\ast}\leq(u_{0})_{\ast}$. But, since, by definition, $(u_{0})_{\ast}\leq u_{0}\leq(u_{0})^{\ast}$ holds, we can conclude that $u_{0}$ is a minimax $L_{H}$-subsolution of (4.1). Again, by Corollary 5.4, it is the only one.
@@
which is absurd.999 Alternatively, one can estimate $H$ by using $z=0$ and Assumption 4.2 (vi) to obtain $\infty\leq C<\infty$ for some constant $C$. Therefore, $(u_{0})^{\ast}<\infty$.

### v2

@@ 5. Comparison principle
(5.1) $\begin{split}&\underline{\mathcal{A}}^{L}w(s_{0},x_{0},\tilde{x}_{0}):=\Big{\{}\varphi\in C^{1,1,1}([s_{0},T]\times\Omega\times\Omega):\,\exists\,T_{0}\in(t_{0},T]:\\ &\qquad 0=(\varphi-w)(s_{0},x_{0},\tilde{x}_{0})=\inf\{(\varphi-w)(t,x,\tilde{x}):\\ &\qquad\qquad\qquad(t,x,\tilde{x})\in[s_{0},T_{0}]\in\mathcal{X}^{L}(s_{0},x_{0})\times\mathcal{X}^{L}(s_{0},\tilde{x}_{0})\Big{\}}.\end{split}$
@@
on $[0,T)\times\Omega\times\Omega$ with parameter $\Upsilon$ if $w$ is u.s.c., $(s,x,\tilde{x})\mapsto\Upsilon(s,x)-w(s,x,\tilde{x})$, $[0,T]\times\Omega\times\Omega\to{\mathbb{R}}$, is l.s.c., and, for every $(s,x,\tilde{x})\in[0,T)\times\Omega\times\Omega$ and every test function $\varphi\in\underline{\mathcal{A}}^{L}w(s,x,\tilde{x})$, we have
@@
Let $L\geq 0$, $u$ be a minimax $L$-subsolution, and $v$ be a minimax $L$-supersolution of (4.1). Then $(t,x,\tilde{x})\mapsto w(t,x,\tilde{x}):=u(t,x)-v(t,\tilde{x})$, $[0,T]\times\Omega\times\Omega\to{\mathbb{R}}$, is a viscosity $L$-subsolution of (5.2) with parameter $\Upsilon=u$.
@@
First note that $u=\Upsilon$ is u.s.c. and $v$ is l.s.c. Thus $w$ is u.s.c. and $(s,x,\tilde{x})\mapsto\Upsilon(s,x)-w(s,x,\tilde{x})=v(s,\tilde{x})$ is l.s.c.
Next, let $(s_{0},x_{0},\tilde{x}_{0})\in[0,T)\times\Omega\times\Omega$, $w(s_{0},x_{0},\tilde{x}_{0})>0$, $\varphi\in\underline{\mathcal{A}}^{L}w(s_{0},x_{0},\tilde{x}_{0})$, and $T_{0}\in(s_{0},T]$ such that (5.1) holds. Let $y_{0}:=u(s_{0},x_{0})$, $\tilde{y}_{0}:=v(s_{0},{\color[rgb]{0,0,0}\tilde{x}_{0}})$, $z:=\partial_{x}\varphi(s_{0},x_{0},\tilde{x}_{0})$, and $\tilde{z}:=-\partial_{\tilde{x}}\varphi(s_{0},x_{0},\tilde{x}_{0})$. By the minimax semisolution properties of $u$ and $v$, there exist $(x,y)\in\mathcal{Y}^{L}(s_{0},x_{0},y_{0},z)$ and $(\tilde{x},\tilde{y})\in\mathcal{Y}^{L}(s_{0},\tilde{x}_{0},\tilde{y}_{0},\tilde{z})$, such that, for every $t\in[s_{0},T]$,
(5.4) $\begin{split}&[u(t,x)-y_{0}]-[v(t,\tilde{x})-\tilde{y}_{0}]\\ &\qquad\geq[y(t)-y_{0}]-[\tilde{y}(t)-\tilde{y}_{0}]\\ &\qquad=\int_{s_{0}}^{t}\left[(x^{\prime}(s),z)-H(s,x,y(s),z)-(\tilde{x}^{\prime}(s),\tilde{z})+H(s,\tilde{x},\tilde{y}(s),\tilde{z})\right]\,{\mathrm{d}}s.\end{split}$
@@
(5.6) $\Psi(s,x,\tilde{x}):=\begin{cases}\frac{{\color[rgb]{0,0,0}(\sup\limits_{t\leq s}\left|{x(t)-\tilde{x}(t)}\right|^{2}-\left|{x(s)-\tilde{x}(s)}\right|^{2})^{2}}}{\sup\limits_{t\leq s}\left|{x(t)-\tilde{x}(t)}\right|^{2}}+\left|{x(s)-\tilde{x}(s)}\right|^{2}&\text{if $\sup\limits_{t\leq s}\left|{x(t)-\tilde{x}(t)}\right|>0$,}\\ 0&\text{otherwise.}\end{cases}$
This functional has been introduced in [zhou2020viscosity]. By Lemma 2.8 in [zhou2020viscosity], it follows that $\Psi\in C^{1,1,1}([s_{0},T]\times\Omega\times\Omega)$ with derivatives $\partial_{t}\Psi(s,x,\tilde{x})=0$ and
(5.7) $\begin{split}&\partial_{x}\Psi(s,x,\tilde{x})=-\partial_{\tilde{x}}\Psi(s,x,\tilde{x})\\ &=\begin{cases}{\tiny\left(2-\frac{{\color[rgb]{0,0,0}4(\sup\limits_{t\leq s}\left|{x(t)-\tilde{x}(t)}\right|^{2}-\left|{x(s)-\tilde{x}(s)}\right|^{2})}}{\sup\limits_{t\leq s}\left|{x(t)-\tilde{x}(t)}\right|^{2}}\right)\cdot[x(s)-\tilde{x}(s)]}&\text{if $\sup\limits_{t\leq s}\left|{x(t)-\tilde{x}(t)}\right|>0$,}\\ 0&\text{otherwise.}\end{cases}\end{split}$
@@
(5.8) $\Phi_{\varepsilon}(s,x,\tilde{x}):=w(s,x,\tilde{x})-\frac{T-s}{T-s_{0}}\cdot\frac{{\color[rgb]{0,0,0}N_{0}}}{2}-\frac{1}{{\varepsilon}}\Psi(s,x,\tilde{x}).$
Fix a point $k_{\varepsilon}=(s_{\varepsilon},x_{\varepsilon},\tilde{x}_{\varepsilon})$ at which the u.s.c. map $\Phi_{\varepsilon}$ attains a maximum on the compact set $[s_{0},T]\times\mathcal{X}^{L}(s_{0},x_{0})\times\mathcal{X}^{L}(s_{0},x_{0})$. Note that
(5.9) ${\color[rgb]{0,0,0}N_{\varepsilon}}:=\Phi_{\varepsilon}(k_{\varepsilon})\geq\Phi_{\varepsilon}(s_{0},x_{0},x_{0})=w(s_{0},x_{0},x_{0})-\frac{{\color[rgb]{0,0,0}N_{0}}}{2}=\frac{{\color[rgb]{0,0,0}N_{0}}}{2}.$
@@
(5.10) $w(k_{\varepsilon})\geq\frac{{\color[rgb]{0,0,0}N_{0}}}{2}+\frac{T-s_{\varepsilon}}{T-s_{0}}\cdot\frac{{\color[rgb]{0,0,0}N_{0}}}{2}+\frac{1}{{\varepsilon}}\Psi(k_{\varepsilon})\geq\frac{{\color[rgb]{0,0,0}N_{0}}}{2}>0.$
Moreover, we have
[…]
and thus $s_{\varepsilon}<T$ if ${\varepsilon}$ is sufficiently small, which we shall assume from now on.
Next, define a map $\varphi_{\varepsilon}:[s_{\varepsilon},T]\times\Omega\times\Omega\to{\mathbb{R}}$ by
$\varphi_{\varepsilon}(s,x,\tilde{x}):={\color[rgb]{0,0,0}N_{\varepsilon}}+\frac{T-s}{T-s_{0}}\cdot\frac{{\color[rgb]{0,0,0}N_{0}}}{2}+\frac{1}{{\varepsilon}}\Psi(s,x,\tilde{x}).$
Then $\varphi_{\varepsilon}\in\underline{\mathcal{A}}^{L}w(k_{\varepsilon})$ with corresponding time $T_{0}=T$ because, by (5.9) and (5.8),
$\varphi_{\varepsilon}(k_{\varepsilon})-w(k_{\varepsilon})$ $={\color[rgb]{0,0,0}N_{\varepsilon}}+\frac{T-s_{\varepsilon}}{T-s_{0}}\cdot\frac{{\color[rgb]{0,0,0}N_{0}}}{2}+\frac{1}{{\varepsilon}}\Psi(k_{\varepsilon})-w(k_{\varepsilon})$
$=w(k_{\varepsilon})-\frac{T-s_{\varepsilon}}{T-s_{0}}\cdot\frac{{\color[rgb]{0,0,0}N_{0}}}{2}-\frac{1}{{\varepsilon}}\Psi(k_{\varepsilon})+\frac{T-s_{\varepsilon}}{T-s_{0}}\cdot\frac{{\color[rgb]{0,0,0}N_{0}}}{2}+\frac{1}{{\varepsilon}}\Psi(k_{\varepsilon})-w(k_{\varepsilon})$
$=0$
$\leq{\color[rgb]{0,0,0}N_{\varepsilon}}-\Phi_{\varepsilon}(s,x,\tilde{x})$
$={\color[rgb]{0,0,0}N_{\varepsilon}}+\frac{T-s}{T-s_{0}}\cdot\frac{{\color[rgb]{0,0,0}N_{0}}}{2}+\frac{1}{{\varepsilon}}\Psi(s,x,\tilde{x})-w(s,x,\tilde{x})$
$=\varphi_{\varepsilon}(s,x,\tilde{x})-w(s,x,\tilde{x})$
@@
$\partial_{t}\varphi_{\varepsilon}(s,x,\tilde{x})=-\frac{{\color[rgb]{0,0,0}N_{0}}}{2(T-s_{0})}\,\,\text{and}\,\,\partial_{x}\varphi_{\varepsilon}(s,x,\tilde{x})=-\partial_{\tilde{x}}\varphi_{\varepsilon}(s,x,\tilde{x})=\frac{1}{{\varepsilon}}\partial_{x}\Psi(s,x,\tilde{x}).$
Hence, since $w$ is a viscosity $L$-subsolution of (5.2) and since (5.10) holds, we have
(5.12) $\begin{split}0&\leq-\frac{{\color[rgb]{0,0,0}N_{0}}}{2(T-s_{0})}+M_{L}\left(1+\left|{\Upsilon(s_{\varepsilon},x_{\varepsilon})}\right|+\frac{\left|{\partial_{x}\Psi(k_{\varepsilon})}\right|}{{\varepsilon}}\right)\cdot\sup_{t\leq s_{\varepsilon}}\left|{x_{\varepsilon}(t)-\tilde{x}_{\varepsilon}(t)}\right|\\ &\leq-\frac{{\color[rgb]{0,0,0}N_{0}}}{2(T-s_{0})}+\tilde{M}_{L}\left(1+\frac{\left|{\partial_{x}\Psi(k_{\varepsilon})}\right|}{{\varepsilon}}\right)\cdot\sup_{t\leq s_{\varepsilon}}\left|{x_{\varepsilon}(t)-\tilde{x}_{\varepsilon}(t)}\right|\end{split}$
for a constant $\tilde{M}_{L}>0$ independent from ${\varepsilon}$. The second line of (5.12) follows from
$\Upsilon(s_{\varepsilon},x_{\varepsilon})$ $\leq\max\{\Upsilon(s,x):(s,x)\in[s_{0},T]\times\mathcal{X}^{L}(s_{0},x_{0})\}<\infty\text{ and}$
$\Upsilon(s_{\varepsilon},x_{\varepsilon})$ $=w(k_{\varepsilon})+[\Upsilon(s_{\varepsilon},x_{\varepsilon})-w(k_{\varepsilon})]$
[…]
$\geq\min\{\Upsilon(s,x)-w(s,x,\tilde{x}):(s,x,\tilde{x})\in[s_{0},T]\times\mathcal{X}^{L}(s_{0},x_{0})\times\mathcal{X}^{L}(s_{0},x_{0})\}$
>$\infty$> - ∞
thanks to $\Upsilon$ being u.s.c., $(s,x,\tilde{x})\mapsto\Upsilon(s,x)-w(s,x,\tilde{x})$ being l.s.c., and $\mathcal{X}^{L}(s_{0},x_{0})$ being compact. It remains to note that, by (5.7),
@@
for every $(s,x,\tilde{x})\in[0,T]\times\Omega\times\Omega$. Together with (5.11), we have
$\left(1+\frac{1}{{\varepsilon}}\left|{\partial_{x}\Psi(k_{\varepsilon})}\right|\right)\cdot\sup_{t\leq s_{\varepsilon}}\left|{x_{\varepsilon}(t)-\tilde{x}_{\varepsilon}(t)}\right|\leq{\color[rgb]{0,0,0}C}\sqrt{\Psi(k_{\varepsilon})}+\frac{C}{{\varepsilon}}\Psi(k_{\varepsilon})\to 0\quad\text{as ${\varepsilon}\downarrow 0$}$
@@
Let $L\geq 0$, $u$ be a minimax $L$-subsolution, and $v$ be a minimax $L$-supersolution of (4.1). Then $u\leq v$ on $[0,T]\times\Omega$.
@@ 6. Existence
We show existence of minimax solutions via Perron’s method (see [BardiCapuzzoDolcetta, section V.2] regarding a corresponding treatment for non-continuous viscosity solutions in the non-path-dependent case). More precisely, the scheme111111 This scheme, which provides existence for minimax solutions, actually predates “Perron’s method” for viscosity solutions introduced in [Ishii87Perron] (we refer to the discussion in [Subbotin, section 10.8] for more details). in [Subbotin, section 8] is adapted to path-dependent Hamilton-Jacobi equations. In the case of Hamiltonians without $u$-dependence, this has already been done in [Lukoyanov03, section 7] (cf. also [BK18JFA, section 5], whose structure we follow here).
@@
where $O_{\delta}(t_{0},x_{0})=\{(t,x)\in[0,T]\times\Omega:\,\mathbf{d}_{\infty}((t_{0},x_{0}),(t_{n},x_{n}))<\delta\}$. We show that $(u_{0})_{\ast}$ is a minimax $L_{H}$-supersolution of (4.1) and $(u_{0})^{\ast}$ is a minimax $L_{H}$-subsolution of (4.1). The comparison principle (Corollary 5.4) yields $(u_{0})^{\ast}\leq(u_{0})_{\ast}$. But, since, by definition, $(u_{0})_{\ast}\leq u_{0}\leq(u_{0})^{\ast}$ holds, we can conclude that $u_{0}$ is a minimax $L_{H}$-solution of (4.1). Again, by Corollary 5.4, it is the only one.
@@
which is absurd.121212 Alternatively, one can estimate $H$ by using $z=0$ and Assumption 4.2 (vi) to obtain $\infty\leq C<\infty$ for some constant $C$. Therefore, $(u_{0})^{\ast}<\infty$.